OpenStudy (anonymous):

If a circle has a radius of 65m, what is the arc length if the central angle is 150°

OpenStudy (anonymous):



OpenStudy (anonymous):

Use this formula for calculating the length of a sector of a circle: \[\frac{ Angle }{ 360 }\times 2\pi r\]

OpenStudy (mayankdevnani):